OpenStudy (ayubie):

I need to find the general solution for: x^2y''+4xy'+2y=2ln(x)+5 I'm not really sure where to start. Do i divide through by the x^2?

OpenStudy (dumbcow):

Yes start by dividing by x^2 Then do a change of variable... \[y = \frac{v}{x^2}\] \[y' = \frac{v'}{x^2} - \frac{2v}{x^3}\] \[y'' = \frac{v''}{x^2} - \frac{4v'}{x^3} +\frac{6v}{x^4}\] After substitution, the equation is... \[\frac{v''}{x^2} = 2 \ln x + 5\] Solve for "v" by integrating twice \[v' = \int\limits v'' dx = \int\limits x^2 (2 \ln x + 5) dx\] \[v = \int\limits v' dx\] ** Use int by parts for above Finally plug in "v" back into ... \[y = \frac{v}{x^2}\] And you have your general solution for y(x) ** Note there will be 2 constants of integration in answer